# Data-Adaptive Identification of Subpopulations Vulnerable to Chemical Exposures using Stochastic Interventions

### v1

@@ 3 Causal Subpopulation Intervention Effects
Our primary objective in this paper is not only to estimate the overall effects of environmental exposures but also to discern how these effects differ across various sub-populations. Within the framework of stochastic shift interventions, our focus is on identifying specific types of individuals from baseline characteristics that change the impact of policy interventions, such as reducing exposure to pollutants. Unlike the previously described stochastic shift parameter, which averages effects across the entire population, we now aim to define our causal parameter for particular subgroups. This allows us to pinpoint the subpopulations where the intervention yields the maximum differential effect, thereby providing more targeted and effective public health strategies.
3.1 Oracle Parameter for Maximal Difference in Expected Outcomes in Subregion
@@
$\psi^{*}=\underset{A_{i}\in\boldsymbol{A},\,\mathcal{V}\subseteq W}{\text{max}}$$\left\{\mathbb{E}\left[Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}-Y\mid W\in\mathcal{V}\right]\right.$ (10)
$\left.-\mathbb{E}\left[Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}-Y\mid W\in\mathcal{V}^{c}\right]\right\},$ (11)
where $A_{i}$ represents an individual exposure within the mixed exposure set $\boldsymbol{A}$. $\mathcal{V}$ and $\mathcal{V}^{c}$ denote subregions of the covariate space $W$. $Y_{A_{\delta_{i}}}$ are the potential outcomes under the specified conditions. The goal is to identify the combination of $A_{i}$ and $\mathcal{V}$ that maximizes the difference in expected outcomes, indicating a modifying effect of the covariate on exposure.
To motivate why we care about the contrast between $W\in\mathcal{V}$ and $W\notin\mathcal{V}$ instead of just looking at the shift effect in the $\mathcal{V}$ region, consider the following. Let $Z$ denote the difference in potential outcomes:
$Z=Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}-Y$
Then, our contrast can be expressed as:
$\mathbb{E}[Z\mid W\in\mathcal{V}]-\mathbb{E}[Z\mid W\notin\mathcal{V}]$
Denote $W\in\mathcal{V}$ as $W=1$ and $W\notin\mathcal{V}$ as $W=0$. By the law of total expectation, the overall average effect $m$ can be expressed as:
$m=\mathbb{E}[Z]=\mathbb{E}[Z\mid W=1]P(W=1)+\mathbb{E}[Z\mid W=0]P(W=0)$
Let $m_{1}=\mathbb{E}[Z\mid W=1]$ and $m_{0}=\mathbb{E}[Z\mid W=0]$. The contrast we care about can be rewritten as:
$m_{1}-m_{0}=\frac{m_{1}-m}{P(W=0)}$
The overall average $m$ is fixed. Therefore, comparing to just $m_{1}$, our contrast will pick regions where $P(W=0)$ is small and $m_{1}$ is large, as opposed to just trying to maximize $m_{1}$. This approach ensures that we do not end up selecting a region that is too small or insignificant in size, which would typically result in a single point, but rather find a meaningful subregion that exhibits a significant differential impact.
In general, we are interested in the causal exposure-covariate region combination where the effect of the intervention, compared to the complementary covariate space, is maximally different, controlling for other exposures. From now on, we refer to these regions as $V$ and $V^{c}$.
@@ 3.2 Statistical Subpopulation Intervention Effects
Here, we explain how to do inference when the data-adaptive parts are fixed, that is, assuming $V,V^{c}$ are fixed. With the causal and statistical parameter determined, we then explain how the regions are generated. Given the same assumptions as stochastic shift interventions, our statistical parameter $\psi$ estimates our causal parameter and is defined as:
@@
$\psi$ $=\underset{A_{i}\in\boldsymbol{A},\,V\subseteq W}{\text{arg max}}\Bigg{(}\mathbb{E}\left[\mathbb{E}\left[Y\mid A_{i}-\delta_{i},\mathbf{A}_{-i},W\in V\right]-\mathbb{E}[Y\mid W\in V]\right]$ (12)
@@
$\quad-\mathbb{E}\left[\mathbb{E}\left[Y\mid A_{i}-\delta_{i},\mathbf{A}_{-i},W\in V^{c}\right]-\mathbb{E}[Y\mid W\in V^{c}]\right]\Bigg{)},$ (13)
This parameter of effect modification, $\Psi_{V}(P)$, is defined as the expected outcome under a stochastic intervention on exposure $A_{i}$, specifically within a subregion $V$ of the covariate space. It integrates over all possible values of $W$ within the subset $V$, capturing the differential impact of $A_{i}$ on the outcome across covariate subgroups.
Operationally, this requires our algorithm to search the exposure-covariate space and, for each exposure, measure the expected outcome change under a shift in regions of the covariates. The algorithm then selects the region and its complementary region where the difference in stochastic intervention impact is maximally different.
The above statistical parameter matches the oracle parameter of interest. However, we may want to include the variance of estimates within each region in order to compare the regions and search for the region difference with the largest statistically significant difference, such as the effect weighed by standard error. To do this, we calculate the influence function in each region, defined below:
@@
$\phi(W,V)=\frac{g(A-\delta|W\in V)}{g(A|W\in V)}\left(Y-\overline{Q}(A,W\in V)\right)+\overline{Q}(A+\delta(W),W\in V)-\Psi(P,V),$ (14)
@@
$\phi(W,V^{c})=\frac{g(A-\delta|W\in V^{c})}{g(A|W\in V^{c})}\left(Y-\overline{Q}(A,W\in V^{c})\right)+\overline{Q}(A+\delta(W),W\in V^{c})-\Psi(P,V^{c}).$ (15)
These influence functions are calculated within the regions $V\in\mathcal{W}$ and $V^{c}\in\mathcal{W}$, allowing us to estimate the effect modification parameter based on a statistical test of differences, like a simple t-statistic.
Although the oracle parameter is defined in terms of population-level effects where everything is perfectly significant, in practical applications, considering statistical significance can provide more robust measures of region selecting in finite samples. Although this approach does not correspond directly to the oracle parameter, it is meaningful to the analyst as it ensures that the identified regions exhibit both substantial and statistically significant differences.
In the method implementation, the package EffectXshift, users can choose whether to find regions with maximum point estimate difference or based on a t-statistic, moving forward we use the t-statistic approach in simulations and applications.
@@ 3.3 Identifying Subpopulations with Maximal Effect Difference
Estimation of the subpopulation intervention effect involves a two-stage procedure. In the first stage we obtain efficient estimation of the individual intervention effects (the expected outcome change for each person, which we denote as IIE) and influence function estimates (IFE) for each exposure and the second stage where we use these estimates in a greedy partitioning algorithm to find the exposure-covariate region with the maximum effect modification.
First Stage: In the first stage, we obtain efficient estimates of the IIE for each exposure $A_{i}$ in $\mathbf{A}$. This is achieved by applying a shift intervention $\delta_{i}$ and debiasing our predictions using TMLE and calculating the expected difference between the outcome under an exposure change compared to no change. Concurrently, we estimate the IFE estimates for each exposure, which quantifies the impact of an individual on our estimator for a shift in the respective exposure.
Specifically, we iterate through the exposures to construct a matrix of IIEs, with each column corresponding to a different exposure. In addition, we create a matrix of IFEs. The first stage consists simply of estimating a stochastic shift intervention for each exposure in the mixture. This is done by 1. Fitting a Super Learner (ensemble of machine learning algorithms that uses cross-validation to select the best estimator or create a linear combination of estimator predictions) for $E[Y|\boldsymbol{A},W]$ which is used to create initial predictions for $\mathbb{E}[Y|A_{i}-\delta,W]$. 2. We then fit a density estimator (or classification estimator - described later) for $g(A_{i}|W)$ and get predictions for $g(A_{i}|W)$ and $g(A_{i}-\delta_{i}|W)$ (or the ratio directly) 3. We then debias our initial predictions for $\mathbb{E}[Y|A_{i}-\delta,W]$ using a least favorable submodel (TMLE) that fits a quasi-binomial regression, regressing the observed outcome (scaled to 0 and 1) onto the clever covariate estimates under no shift offset by the initial predictions from the Super Learner. Using this model, we then input the initial predictions given a shift and density given a shift and get updated predictions through the model which are debiased towards our shift parameter of interest. We then plug these estimates in our influence function to get the ICE for each observation.
We loop through each exposure in the mixture and get debiased estimates of the effect of a shift $\delta_{i}$ in each exposure $A_{i}$, constructing a matrix of individual intervention effects. We also construct a matrix for the influence curve estimates for each intervention in each exposure.
Second Stage: In the second stage, we loop through the IIE and IC estimates for each exposure and regress these onto the covariates using a greedy partitioning algorithm. This algorithm searches the covariate space and finds the region and it’s complement which has the maximum effect difference based on a t-statistic, comparing the averages from the IIE and standard error estimates from the IC in each covariate region and splitting based on decreasing p-value. The detailed pseudocode in Algorithm 1 describes the partitioning approach, using the t-statistic to find the exposure-covariate region that has the maximum difference in subpopulation intervention effect.
@@
1:Input: Data $D$, covariates $W$, individual intervention effect $IIE$, influence curve estimates $ICE$, max depth $d_{\text{max}}$, minimum observations $min\_obs$
2:Output: Rule that demarcates the covariate region with the expected outcome difference
3:procedure T-Part($D$, $W$, $IIE$, $ICE$, depth=0)
@@
5:         return $best\_split$triangleright Stop recursion
@@
7:     Init $best\_split\leftarrow\text{Null}$, $p_{\text{best}}\leftarrow 1$
8:     for $w\in W$ do
9:         for $s$ in unique values of $w$ do
10:              $\mathcal{W}^{*}\leftarrow(D[w]\leq s)$
11:              if sum($\mathcal{W}^{*}$) ¡ $min\_obs$ or sum($\neg\mathcal{W}^{*}$) ¡ $min\_obs$ then
@@
14:              Split $D$ into $D_{\text{left}}$ and $D_{\text{right}}$ based on $\mathcal{W}^{*}$
15:              Calculate mean IIE for $D_{\text{left}}$ and $D_{\text{right}}$
16:              Calculate variance of ICE for $D_{\text{left}}$ and $D_{\text{right}}$
17:              Calculate t-statistic using:
@@
18:              Calculate p-value from t-stat
19:              if p-value $<p_{\text{best}}$ then
20:                  $p_{\text{best}}\leftarrow p\_value$
21:                  $best\_split\leftarrow\{w,s,p_{\text{best}}\}$
22:              end if
23:         end for
24:     end for
25:     if $best\_split\neq\text{Null}$ then
26:         Split $D$ into $D_{\text{left}}$ and $D_{\text{right}}$ based on $best\_split$
27:         Recursively apply T-Part($D_{\text{left}},W,IIE,ICE,depth+1$)
28:         Recursively apply T-Part($D_{\text{right}},W,IIE,ICE,depth+1$)
29:     else
30:         return No more significant splits found $\triangleright$ Stop recursion
31:     end if
32:end procedure
33:Main Loop:
34:for $IIE_{i}\in 1:\text{ncol}(IIE)$ do
35:     Initiate T-Part($D$, $W$, $IIE_{i}$, $ICE_{i}$, 0)
36:end for
37:Select the top max effect modifier-covariate relationship: extract region rule
38:Return $max\_mod\_rule$

### v2

@@ 3 Causal Subpopulation Intervention Effects
Our primary objective in this paper is not only to estimate the overall effects of environmental exposures but also to discern how these effects differ across various subpopulations. Within the framework of stochastic shift interventions, our focus is on identifying specific types of individuals from baseline characteristics that change the impact of policy interventions, such as reducing exposure to pollutants. Unlike the previously described stochastic shift parameter, which averages effects across the entire population, we aim to define our causal parameter for particular subgroups. This allows us to pinpoint the subpopulations where the intervention yields the maximum differential effect, thereby providing more targeted and effective public health strategies.
3.1 Oracle Parameter for Maximal Difference in Expected Outcomes in Subregions
@@
$\psi^{*}$ $=\underset{A_{i}\in\boldsymbol{A},\,\mathcal{V}\subseteq\mathcal{W}}{\sup}\left\{\mathbb{E}\left[Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}-Y\mid W\in\mathcal{V}\right]\right.$
$\left.\quad-\mathbb{E}\left[Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}-Y\mid W\in\mathcal{V}^{c}\right]\right\},$ (3)
where:
  • $\boldsymbol{A}=(A_{1},A_{2},\ldots,A_{p})$ denotes the vector of exposures in the mixture, with $A_{i}$ representing the $i$-th exposure and $\mathbf{A}_{-i}$ denoting all exposures excluding $A_{i}$.
  • $\mathcal{V}$ and $\mathcal{V}^{c}$ denote a subregion of the covariate space $\mathcal{W}$ and its complement, respectively.
  • $Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}$ represents the potential outcome under a stochastic shift intervention where exposure $A_{i}$ is shifted by $\delta_{i}$, while other exposures $\mathbf{A}_{-i}$ remain unchanged.
$\psi^{*}$ identifies the combination of exposure $A_{i}$ and subregion $\mathcal{V}$ that maximizes the difference in expected outcomes between $\mathcal{V}$ and its complement $\mathcal{V}^{c}$. This parameter captures effect modification, highlighting where the intervention has the most substantial and relevant impact.
In general, we are interested in the causal exposure-covariate region combination where the effect of the intervention, compared to the complementary covariate space, is maximally different, controlling for other exposures. From now on, we refer to these regions as $\mathcal{V}$ and $\mathcal{V}^{c}$.
@@ 3.2 Statistical Subpopulation Intervention Effects
This section reformulates the causal parameter of interest in terms of observable quantities, thereby defining a statistical parameter that can be estimated from the data. We consider a fixed subpopulation $\mathcal{V}\subseteq\mathcal{W}$, but emphasize that subpopulation-specific parameters are interpreted as features of the full population distribution. Let $\pi_{\mathcal{V}}=P(W\in\mathcal{V})$ denote the probability that a randomly selected unit from the full population belongs to $\mathcal{V}$.
Under standard identification assumptions, including no unmeasured confounding and positivity, the causal parameter $\psi^{*}$ is identified as:
@@
$\psi^{*}$ $=\sup_{A_{i}\in\boldsymbol{A},\;\mathcal{V}\subseteq\mathcal{W}}\left\{\mathbb{E}\bigl{[}\overline{Q}_{0}(A_{i}-\delta_{i},\mathbf{A}_{-i},W)-\overline{Q}_{0}(A_{i},\mathbf{A}_{-i},W)\mid W\in\mathcal{V}\bigr{]}\right.$
@@
$\left.\quad-\mathbb{E}\bigl{[}\overline{Q}_{0}(A_{i}-\delta_{i},\mathbf{A}_{-i},W)-\overline{Q}_{0}(A_{i},\mathbf{A}_{-i},W)\mid W\in\mathcal{V}^{c}\bigr{]}\right\},$ (4)
where $\overline{Q}_{0}(a_{i},\mathbf{a}_{-i},w)=\mathbb{E}[Y\mid A_{i}=a_{i},\mathbf{A}_{-i}=\mathbf{a}_{-i},W=w]$.
Fixing $\mathcal{V}$, we define the statistical parameter $\psi_{\mathcal{V}}$ as:
$\psi_{\mathcal{V}}$ $=\max_{A_{i}\in\boldsymbol{A}}\bigl{(}\Psi_{\mathcal{V}}(P)-\Psi_{\mathcal{V}^{c}}(P)\bigr{)},$ (5)
where
$\Psi_{\mathcal{V}}(P)$ $=\frac{\mathbb{E}[\Delta_{i}(W)\mathbb{I}\{W\in\mathcal{V}\}]}{\pi_{\mathcal{V}}},\quad\Psi_{\mathcal{V}^{c}}(P)=\frac{\mathbb{E}[\Delta_{i}(W)\mathbb{I}\{W\in\mathcal{V}^{c}\}]}{1-\pi_{\mathcal{V}}},$ (6)
$\Delta_{i}(W)$ $=\overline{Q}_{0}(A_{i}-\delta_{i},\mathbf{A}_{-i},W)-\overline{Q}_{0}(A_{i},\mathbf{A}_{-i},W).$ (7)
This formulation ensures that $\Psi_{\mathcal{V}}(P)$ and $\Psi_{\mathcal{V}^{c}}(P)$ are understood as expectations taken over the full population distribution, normalized by $\pi_{\mathcal{V}}$ and $1-\pi_{\mathcal{V}}$ respectively. The parameter $\psi_{\mathcal{V}}$ thus represents a feature of the entire population, focusing attention on differences between a particular subpopulation $\mathcal{V}$ and its complement.
Influence Functions Corresponding to the Parameters
To perform inference on $\Psi_{\mathcal{V}}(P)$ and $\Psi_{\mathcal{V}^{c}}(P)$, we rely on their efficient influence functions (EIFs), which are essential for constructing asymptotically efficient estimators and valid confidence intervals. The EIFs incorporate the normalization by $\pi_{\mathcal{V}}$:
@@
$\phi_{\mathcal{V}}(O)$ $=\frac{\mathbb{I}\{W\in\mathcal{V}\}}{\pi_{\mathcal{V}}}\left[\frac{g_{0}(A_{i}-\delta_{i}\mid W)}{g_{0}(A_{i}\mid W)}\bigl{(}Y-\overline{Q}_{0}(A_{i},\mathbf{A}_{-i},W)\bigr{)}+\Delta_{i}(W)-\Psi_{\mathcal{V}}(P)\right]$
$\quad+\bigl{(}\mathbb{I}\{W\in\mathcal{V}\}-\pi_{\mathcal{V}}\bigr{)}\frac{\Psi_{\mathcal{V}}(P)}{\pi_{\mathcal{V}}},$ (8)
@@
$\phi_{\mathcal{V}^{c}}(O)$ $=\frac{\mathbb{I}\{W\in\mathcal{V}^{c}\}}{1-\pi_{\mathcal{V}}}\left[\frac{g_{0}(A_{i}-\delta_{i}\mid W)}{g_{0}(A_{i}\mid W)}\bigl{(}Y-\overline{Q}_{0}(A_{i},\mathbf{A}_{-i},W)\bigr{)}+\Delta_{i}(W)-\Psi_{\mathcal{V}^{c}}(P)\right]$
$\quad+\bigl{(}\mathbb{I}\{W\in\mathcal{V}^{c}\}-(1-\pi_{\mathcal{V}})\bigr{)}\frac{\Psi_{\mathcal{V}^{c}}(P)}{1-\pi_{\mathcal{V}}}.$ (9)
Estimation and Inference
In practice, the true functions $\overline{Q}_{0}$ and $g_{0}$, as well as $\pi_{\mathcal{V}}$, are unknown and must be estimated from observed data. Let $\hat{\pi}_{\mathcal{V}}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\{W_{i}\in\mathcal{V}\}$. We define $\hat{g}(A\mid W)$ as a suitable estimator of $g_{0}(A\mid W)$, the conditional density of $A$ given $W$. The function $\overline{Q}_{0}(A,W)$ is estimated by $\hat{\overline{Q}}(A,W)$, and we employ Targeted Maximum Likelihood Estimation (TMLE) to refine this initial estimate, resulting in $\hat{\overline{Q}}^{\ast}(A,W)$. The superscript $\ast$ indicates that the estimate has been updated in a manner that targets the parameter of interest, thereby reducing bias and improving efficiency.
Once $\hat{\overline{Q}}^{\ast}$ and $\hat{g}$ are obtained, we define:
$\hat{\Delta}_{i}(W_{i})$ $=\hat{\overline{Q}}^{\ast}(A_{i}-\delta_{i},\mathbf{A}_{-i},W_{i})-\hat{\overline{Q}}^{\ast}(A_{i},\mathbf{A}_{-i},W_{i}).$ (10)
We then estimate $\Psi_{\mathcal{V}}(P)$ and $\Psi_{\mathcal{V}^{c}}(P)$ as:
$\hat{\Psi}_{\mathcal{V}}$ $=\frac{\sum_{i=1}^{n}\hat{\Delta}_{i}(W_{i})\mathbb{I}\{W_{i}\in\mathcal{V}\}}{\hat{\pi}_{\mathcal{V}}},\quad\hat{\Psi}_{\mathcal{V}^{c}}=\frac{\sum_{i=1}^{n}\hat{\Delta}_{i}(W_{i})\mathbb{I}\{W_{i}\in\mathcal{V}^{c}\}}{1-\hat{\pi}_{\mathcal{V}}}.$ (11)
Let $n_{\mathcal{V}}=\sum_{i=1}^{n}\mathbb{I}\{W_{i}\in\mathcal{V}\}$ and $n_{\mathcal{V}^{c}}=\sum_{i=1}^{n}\mathbb{I}\{W_{i}\in\mathcal{V}^{c}\}$. The parameter of interest, representing the difference in subpopulation-specific effects, is estimated by:
$\hat{\psi}_{\mathcal{V}}=\hat{\Psi}_{\mathcal{V}}-\hat{\Psi}_{\mathcal{V}^{c}}.$
Variance Estimation and Statistical Inference
To derive standard errors and confidence intervals, we use the estimated influence functions:
$\hat{\phi}_{\mathcal{V}}(O_{i}),\quad\hat{\phi}_{\mathcal{V}^{c}}(O_{i}),$
obtained by plugging the estimated quantities $\hat{\overline{Q}}^{\ast}$, $\hat{g}$, and $\hat{\pi}_{\mathcal{V}}$ into the EIF formulas.
The variance of $\hat{\psi}_{\mathcal{V}}$ is:
$\widehat{\mathrm{Var}}(\hat{\psi}_{\mathcal{V}})=\widehat{\mathrm{Var}}(\hat{\Psi}_{\mathcal{V}})+\widehat{\mathrm{Var}}(\hat{\Psi}_{\mathcal{V}^{c}}),$
assuming negligible covariance due to disjoint subpopulations. The variance estimators require incorporating the appropriate scaling by $1/n_{\mathcal{V}}$ and $1/n_{\mathcal{V}^{c}}$. After computing these variances from the EIFs, a normal approximation can be employed to construct confidence intervals and conduct hypothesis tests, such as testing whether $\psi_{\mathcal{V}}=0$.
In summary, defining the parameter $\psi_{\mathcal{V}}$ as a population-level feature involving a particular subpopulation $\mathcal{V}$ and using the normalized EIFs ensures valid statistical inference. By applying TMLE and variance estimation through the EIFs, this framework provides efficient, assumption-lean estimates for subpopulation-specific intervention effects and the differences between subpopulations. Such estimates facilitate informed decision-making and targeted interventions in settings where heterogeneous effects are of interest.
Incorporating Variability in Region Selection
Although the statistical parameter $\psi_{\mathcal{V}}$ quantifies the difference in expected outcomes between the subpopulation $\mathcal{V}$ and its complement $\mathcal{V}^{c}$, it is essential to consider the variability of these estimates. Even a large observed difference may not be statistically significant once variability is taken into account. By leveraging the influence functions and their corresponding variance estimates, we can formally assess the statistical significance of the observed effect modification.
To test the null hypothesis $H_{0}:\psi_{\mathcal{V}}=0$, we construct a t-statistic:
$T_{\mathcal{V}}=\frac{\hat{\psi}_{\mathcal{V}}}{\sqrt{\widehat{\mathrm{Var}}(\hat{\psi}_{\mathcal{V}})}}.$
Under standard regularity conditions, $T_{\mathcal{V}}$ converges in distribution to a standard normal random variable. Consequently, we can employ conventional inferential procedures (e.g., using standard normal quantiles) to determine whether the observed difference in intervention effects is statistically significant. If $|T_{\mathcal{V}}|$ exceeds a chosen critical value (e.g., corresponding to a two-sided $\alpha=0.05$ level), we reject $H_{0}$ and conclude that there is evidence of a meaningful difference in the intervention effects between $\mathcal{V}$ and $\mathcal{V}^{c}$.
Remark
In applied settings, it is crucial to consider both the magnitude and statistical significance of the estimated effect difference. Focusing solely on the point estimate without accounting for variability risks identifying subpopulations based on random fluctuations rather than true differences. By incorporating a careful variance-based assessment, researchers can more confidently identify meaningful subpopulations for targeted interventions and informed policy decisions.
@@ 3.3 Identifying Subpopulations with Maximal Effect Difference
Building upon the statistical parameters and influence functions defined in the previous section, we now describe a procedure for identifying subpopulations with maximal effect modification for a given exposure. The estimation involves a two-stage approach:
First Stage: We obtain efficient estimates of the Individual Intervention Effects (IIEs) and the corresponding Influence Curve Estimates (ICEs) for each exposure $A_{k}$ in the mixture $\boldsymbol{A}=(A_{1},A_{2},\ldots,A_{p})$.
Second Stage: Utilizing these estimates, we apply a greedy partitioning algorithm to find covariate regions $\mathcal{V}$ where the effect modification $\psi_{\mathcal{V}}$ is maximized and statistically significant.
First Stage: Estimating IIEs and ICEs
For each exposure $A_{k}$, we proceed as follows:
  1. 1.
    Initial Estimation of Outcome Regression: Fit an initial estimator $\hat{\overline{Q}}(A,W)$ for $\mathbb{E}[Y\mid\boldsymbol{A},W]$ using flexible machine learning methods (e.g., Super Learner).
  2. 2.
    Estimating the Density Ratio: Estimate $\hat{g}(A_{k}\mid W)$ and $\hat{g}(A_{k}-\delta_{k}\mid W)$, or directly estimate the density ratio $\hat{r}(A_{k},W)=\frac{\hat{g}(A_{k}-\delta_{k}\mid W)}{\hat{g}(A_{k}\mid W)}$.
  3. 3.
    Debiasing with TMLE: Update the initial estimator $\hat{\overline{Q}}(A,W)$ to $\hat{\overline{Q}}^{\ast}(A,W)$ via Targeted Maximum Likelihood Estimation (TMLE), targeting the parameter $\Psi_{\mathcal{V}}(P)$ for the chosen exposure. This involves:
    1. (a)
      Constructing the clever covariate $H(A_{k},W)=\hat{r}(A_{k},W)$.
    2. (b)
      Fitting a parametric model (e.g., logistic regression for binary outcomes) of $Y$ on $H(A_{k},W)$ with offset $\text{logit}(\hat{\overline{Q}}(A_{k},W))$ for binary $Y$, or a suitable quasi-likelihood model for continuous $Y$.
    3. (c)
      Obtaining $\hat{\overline{Q}}^{\ast}(A,W)$, the updated estimator incorporating the TMLE fluctuation.
  4. 4.
    Calculating Individual-Level Estimates: For each individual $i$:
    1. (a)
      Individual Intervention Effect (IIE): Compute the estimated IIE for exposure $A_{k}$:
      $\hat{\Delta}_{ik}(W_{i})=\hat{\overline{Q}}^{\ast}(A_{k}-\delta_{k},\,\mathbf{A}_{-k},\,W_{i})-\hat{\overline{Q}}^{\ast}(A_{k},\,\mathbf{A}_{-k},\,W_{i}).$ (12)
    2. (b)
      Influence Curve Estimate (ICE): Compute the estimated influence function:
      $\hat{\phi}_{ik}(O_{i})=\frac{\hat{g}(A_{k}-\delta_{k}\mid W_{i})}{\hat{g}(A_{k}\mid W_{i})}\bigl{(}Y_{i}-\hat{\overline{Q}}^{\ast}(A_{k},W_{i})\bigr{)}+\hat{\Delta}_{ik}(W_{i})-\hat{\Psi}_{k},$ (13)
      where $\hat{\Psi}_{k}$ is the estimated population-level parameter (e.g., the average intervention effect for exposure $A_{k}$ in the sample).
By iterating through each exposure $A_{k}$, we construct matrices of IIEs $\hat{\Delta}$ and ICEs $\hat{\phi}$, where rows correspond to individuals $i$ and columns correspond to exposures $k$.
Second Stage: Identifying Covariate Regions with Maximal Effect Modification
In the second stage, we utilize the estimated IIEs $\hat{\Delta}_{ik}$ and ICEs $\hat{\phi}_{ik}$ for a chosen exposure $A_{k}$ to identify covariate regions $\mathcal{V}$ where the effect modification $\psi_{\mathcal{V}}$ is maximized. We employ a greedy partitioning algorithm based on statistical significance, ensuring that the identified regions exhibit both substantial and statistically significant differences.
The algorithm proceeds as follows:
  1. 1.
    Initialization: Start with the full covariate space $\mathcal{W}$.
  2. 2.
    Greedy Search for Optimal Split: For each covariate $W_{j}$ and each candidate split point $s$:
    1. (a)
      Define $\mathcal{V}=\{i:W_{ij}\leq s\}$ and $\mathcal{V}^{c}=\{i:W_{ij}>s\}$.
    2. (b)
      Compute the estimated effect modification:
      $\hat{\psi}_{\mathcal{V}}=\hat{\Psi}_{\mathcal{V}}-\hat{\Psi}_{\mathcal{V}^{c}},$
      where
      $\hat{\Psi}_{\mathcal{V}}=\frac{1}{n_{\mathcal{V}}}\sum_{i\in\mathcal{V}}\hat{\Delta}_{ik}(W_{i}),\quad\hat{\Psi}_{\mathcal{V}^{c}}=\frac{1}{n_{\mathcal{V}^{c}}}\sum_{i\in\mathcal{V}^{c}}\hat{\Delta}_{ik}(W_{i}).$
    3. (c)
      Estimate the variance $\widehat{\mathrm{Var}}(\hat{\psi}_{\mathcal{V}})$ using the ICEs:
      $\widehat{\mathrm{Var}}(\hat{\psi}_{\mathcal{V}})=\frac{1}{n_{\mathcal{V}}^{2}}\sum_{i\in\mathcal{V}}\bigl{(}\hat{\phi}_{ik}(O_{i})\bigr{)}^{2}\;+\;\frac{1}{n_{\mathcal{V}^{c}}^{2}}\sum_{i\in\mathcal{V}^{c}}\bigl{(}\hat{\phi}_{ik}(O_{i})\bigr{)}^{2}.$
    4. (d)
      Compute the t-statistic:
      $T_{\mathcal{V}}=\frac{\hat{\psi}_{\mathcal{V}}}{\sqrt{\widehat{\mathrm{Var}}(\hat{\psi}_{\mathcal{V}})}}.$
    5. (e)
      Record the split that maximizes $|T_{\mathcal{V}}|$, subject to minimum sample size constraints to ensure valid asymptotic approximations.
  3. 3.
    Recursive Partitioning: Recursively apply the above procedure to the resulting subpopulations, increasing the depth of the partitioning tree until a predefined maximum depth is reached or no further statistically significant splits are found.
  4. 4.
    Selection of Optimal Region: Among all considered splits, select the covariate $W_{j}$ and split point $s$ that yield the maximal statistically significant effect modification.
Algorithm Implementation
The detailed pseudocode in Algorithm 1 outlines the recursive partitioning approach, leveraging the t-statistic to identify the covariate regions that maximize the difference in subpopulation intervention effects.
@@
1:Input: Data $D$, covariates $W$, IIEs $\hat{\Delta}$, ICEs $\hat{\phi}$, maximum depth $d_{\text{max}}$, minimum observations min_obs
2:Output: Rule defining the covariate region $\mathcal{V}$ with maximal effect modification
3:procedure T-Part($D$, $W$, $\hat{\Delta}$, $\hat{\phi}$, depth=0)
@@
5:         return $\triangleright$ Maximum depth reached
@@
[…]
8:     for each covariate $W_{j}$ do
9:         for each possible split point $s$ in $W_{j}$ do
10:              Define $\mathcal{V}=\{i:W_{ij}\leq s\}$, $\mathcal{V}^{c}=\{i:W_{ij}>s\}$
[…]
@@
14:              Compute $\hat{\psi}_{\mathcal{V}}$ and $\widehat{\mathrm{Var}}(\hat{\psi}_{\mathcal{V}})$ as above
15:              Compute $T_{\mathcal{V}}=\dfrac{\hat{\psi}_{\mathcal{V}}}{\sqrt{\widehat{\mathrm{Var}}(\hat{\psi}_{\mathcal{V}})}}$
16:              if  $|T_{\mathcal{V}}|>|T_{\text{best}}|$  then
17:                  $T_{\text{best}}\leftarrow T_{\mathcal{V}}$
@@
19:              end if
20:         end for
21:     end for
[…]
23:         Partition data based on best_split into $D_{\mathcal{V}}$ and $D_{\mathcal{V}^{c}}$
24:         Recursively apply T-Part($D_{\mathcal{V}},W,\hat{\Delta},\hat{\phi},\text{depth}+1$)
25:         Recursively apply T-Part($D_{\mathcal{V}^{c}},W,\hat{\Delta},\hat{\phi},\text{depth}+1$)
26:     else
27:         return $\triangleright$ No significant split found
28:     end if
29:end procedure
30:Main Loop:
31:for each exposure $A_{i}$ in $\boldsymbol{A}$ do
32:     Initiate T-Part($D$, $W$, $\hat{\Delta}_{i}$, $\hat{\phi}_{i}$, 0)
33:end for
34:Select the covariate region $\mathcal{V}$ with the maximal significant effect modification
35:Return $\mathcal{V}$ and the corresponding effect modification estimate $\hat{\psi}_{\mathcal{V}}$
Remark. By integrating the estimated influence functions into the partitioning algorithm, we ensure that the selection of covariate regions is guided not only by the magnitude of the estimated effect differences but also by their statistical significance. This approach helps to mitigate the risk of overfitting and the identification of spurious subpopulations due to random variability.